OpenStudy (anonymous):

Maggie has eaten 600 calories today. Her nutritionist instructed her to eat 1,500 calories per day. Of those 1,500 calories, at least 29% are supposed to come from protein. Of her 600 calories consumed today, 24% have been from protein. How many more calories (C) from protein does Maggie need to eat today in order to have at least 29%? 291 < C < 900 174 < C < 435 435 < C < 900 144 < C < 435

OpenStudy (anonymous):

1500*0.29=435 how many calories have to come from protein total 600*0.24=144 how many calories from protein she has eaten today 435-144=291 how many calories from protein she still needs to eat